OpenStudy (anonymous):

Find x and y if: (5y+3xi)+(2x- 7yi)=11-27i How do I solve this?

OpenStudy (anonymous):

the real terms are equal and the imaginary terms are equal 5y+2x=11 and 3x-7y=-27

OpenStudy (anonymous):

Yup I got that far, I'm just stuck after that.

OpenStudy (anonymous):

6x+15y=33 -6x+14y=54 29y=87 y=3 substitute back and get 5(3)+2x=11 2x=-4 x=-2 check (5(3)+3(-2)i)+(2(-2)- 7(3)i)=11-27i 15-6i+-4-21i=11-27i 11-27i=11-27i

OpenStudy (anonymous):

Where does the 29y=87 come from? What was each term multiplied by?

OpenStudy (anonymous):

add those two equations together

OpenStudy (anonymous):

I see that, but wouldn't you have had to alter the original equations to get the both to equal 6 and -6?

OpenStudy (anonymous):

Ohhhh! Nevrmind, I had the first equation written as 5x+2y=11:) Sorry

OpenStudy (anonymous):

so multiple it by 2 and the other one by 3

OpenStudy (anonymous):

no problem. glad you understand. let me know if anything else seems confusing

OpenStudy (anonymous):

Thank you sooo much mate!!